# LLaRA: Supercharging Robot Learning Data for Vision-Language Policy

### v2

@@ Abstract
LLMs with visual inputs, i.e., Vision Language Models (VLMs), have the capacity to process state information as visual-textual prompts and respond with policy decisions in text. We propose LLaRA: Large Language and Robotics Assistant, a framework that formulates robot action policy as conversations and provides improved action outputs when trained with auxiliary data that complements policy learning. We first introduce an automated pipeline to generate conversation-style instruction tuning data from existing behavior cloning data. Then we enrich the dataset in a self-supervised fashion by formulating six auxiliary tasks. A VLM finetuned with the resulting collection of datasets can generate meaningful robot action policy decisions. Our experiments across multiple simulated and real-world environments demonstrate the state-of-the-art performance of the proposed LLaRA framework. The code, datasets, and pretrained models are available at https://github.com/LostXine/LLaRA.
Figure 1: A real-world demonstration of LLaRA solving an unseen task. In the training stage, expert trajectories are converted into instruction-tuning data with image coordinates. Additionally, auxiliary data is generated to enhance the policy in a self-supervised manner. Using this data, we finetune a pretrained Vision Language Model (VLM) as a robot policy. At inference, the VLM is queried to generate actions in natural language.
@@ 1 Introduction
Several robot learning approaches also leverage pretrained LLMs / VLMs (Zeng et al., 2022; Brohan et al., 2023c; Driess et al., 2023; Vemprala et al., 2024; Kim et al., 2024; Niu et al., 2024) with recent work focusing on low-level robot actions (Brohan et al., 2023b; a; Padalkar et al., 2023). The strong performance of such approaches can be attributed to the extensive world knowledge (e.g., understanding of physics, human common-sense) (Yu et al., 2023; Zhao et al., 2024) and powerful reasoning abilities (Creswell & Shanahan, 2022; Liu et al., 2023b) of underlying language models. However, studies on curating and using conversation-style data for instruction tuning in robotics have been rather limited (Zhen et al., 2024).
Motivated by the promising attributes of VLMs, we explore a formulation of VLM-based robot action policy in this paper. Our goal is to adapt a VLM as a robot action policy that can handle diverse visuomotor control challenges, a process we call Visuomotor Instruction Tuning. More specifically, we transform a typical behavior cloning (BC) dataset into an instruction dataset, Instruct-BC (inBC), which is then used to finetune a VLM. Given a state described in visual-textual modalities, our VLM is trained to generate suitable actions as text. Such a formulation based on conversation-style instruction-response data enables us to instruction-tune and convert a VLM into a robot action policy effortlessly. However, we also note that the effectiveness of an instruction-tuned VLM depends heavily on the quality of data formulation (Liu et al., 2023a; Ranasinghe et al., 2024b; Kuckreja et al., 2024), which is non-trivial to automate (or, scale) for new domains  (Li et al., 2023; Kuckreja et al., 2024; Hong et al., 2023; Chen et al., 2023; Ranasinghe et al., 2024b). Therefore, to strengthen (or, supercharge) such data, we construct auxiliary datasets that complement policy learning from the same BC dataset, in a self-supervised fashion without any extra action label.
Our overall framework termed LLaRA (Large Language and Robotics Assistant) generates visuomotor instruction data that can efficiently and effectively finetune a VLM into a robot action policy. It is motivated by LLaVA (Liu et al., 2023a) which is designed primarily for vision tasks. Similarly, LLaRA offers the entire framework of data generation, model formulation, and training pipeline for VLMs, now specialized for robot learning. Our key contributions are:
@@
    Formulating robot manipulation tasks into instruction-response pairs described by natural language, which enables successful instruction tuning of a VLM as a policy.
@@
    Aligning robot actions to image coordinates for VLMs, making the transfer of a pretrained VLM to the robotics domain easier, particularly when the training data is limited.
@@
    Identifying and generating auxiliary instruction data that further enhances robot policy learning in a self-supervised manner.
@@ 2 Related work
PALM-E (Driess et al., 2023) utilizes a pretrained VLM for embodied reasoning and planning. However, it assumes access to additional policies capable of performing low-level skills from a limited vocabulary, with the VLM generating natural language that conditions these low-level commands. Our method requires neither a language-conditioned low-level controller, which can be expensive to learn for each environment, nor a lot of training data to achieve good performance. In addition, because we use 2D image coordinates to present position, one can easily transfer our model to another embodiment/domain with limited training samples. SpatialVLM (Chen et al., 2024) focuses on enhancing the spatial reasoning ability of the VLM, resulting in a VLM that serves as a spatial relationship predictor rather than a robot policy. To embed SpatialVLM into a policy, chain-of-thought spatial reasoning is required. Our method explicitly transforms a pretrained VLM into a robot policy that directly generates robot actions, resulting in a simpler framework and no need for an additional policy. In addition, SpatialVLM studies only the spatial relationships between objects. Our method also learns temporal environment dynamics through action prediction and future prediction datasets, making our VLM better prepared as a robot policy.
@@
There are concurrent works (Kim et al., 2024; Yuan et al., 2024; Niu et al., 2024) which also utilize a LLaVA style VLM as a robot policy. All such approaches employ a VLM as a robot action policy, processing both visual observations and textual instructions as inputs. OpenVLA (Kim et al., 2024) employs specialized tokens to represent quantized actions, a method akin to RT-2 (Brohan et al., 2023a). OpenVLA also incorporates DINOv2 (Oquab et al., 2023) as an additional visual encoder along with SigLIP (Zhai et al., 2023), compared to a single CLIP (Radford et al., 2021) vision encoder used in LLaVA. RoboPoint (Yuan et al., 2024) introduces a novel point-based action space and a scalable data pipeline tailored for spatial affordance prediction. Actions are delineated via points on an RGB image, which are subsequently translated to 3D space using depth data. This approach eliminates the reliance on predefined action primitives (Liang et al., 2023; Singh et al., 2023), external object detectors (Huang et al., 2023b; Liu et al., 2024a), and iterative visual prompting (Nasiriany et al., 2024). Another concurrent work, LLARVA (Niu et al., 2024), differs by generating both 2D visual traces in image coordinates and corresponding textual actions as outputs, with the former functioning as an auxiliary task. In contrast, our method describes robot action using image coordinates in natural language, which creates a more intuitive link between vision and action without modifying the existing VLM model architecture. Moreover, all the aforementioned studies lack of comprehensive investigation into the methods for generating auxiliary datasets from existing robot data, as well as the implications of integrating such datasets. Our work LLaRA distinctively addresses this gap, marking a critical divergence from the concurrent studies. That is, in addition to directly supervising a VLM based on robot actions, the LLaRA framework automatically generates multiple auxiliary datasets with complementary training objectives from the existing robot trajectories. Such auxiliary datasets are designed to train a VLM to capture spatial/temporal relations between objects in the scene, enabling better scene understanding. We experimentally confirm that such additional supervision significantly benefits robot action policy learning.
@@ 3 Preliminary: visual instruction tuning
Visual Instruction Tuning is a framework to finetune a large Vision Language Model (VLM) using task-specific instructions. The objective is to develop a general-purpose multimodal model capable of solving diverse vision tasks described by language (Huang et al., 2023a). A typical example of Visual Instruction Tuning is LLaVA (Liu et al., 2023a), which consists of three neural networks as its model: a large language model $\theta_{\text{LLM}}$, a visual encoder $\theta_{V}$ and an adapter layer $\theta_{\text{MLP}}$ (see Fig. 2).
Consider a conversation data sample that contains a single image $x_{v}$ and user instruction text $x_{l}$ as inputs. During inference, the visual encoder $\theta_{V}$ and adapter layer $\theta_{\text{MLP}}$ process $x_{v}$ successively to produce a set of visual tokens that can be directly concatenated with the textual tokens of $x_{l}$ in language embedding space of LLM $\theta_{\text{LLM}}$. Next, $\theta_{\text{LLM}}$ autoregressively generates new tokens conditioned on both the visual tokens from $x_{v}$ and text tokens from $x_{l}$. These new tokens will be decoded into natural language as the output of the model.
A two-phase training process is executed in Liu et al. (2023a) with a next-token prediction objective over the $\theta_{\text{LLM}}$ output in each phase. First, only $\theta_{\text{MLP}}$ is trained using image-caption pairs (similar to datasets for CLIP (Radford et al., 2021) training). Then, both $\theta_{\text{MLP}}$ and $\theta_{\text{LLM}}$ are finetuned using a specialized instruction tuning dataset termed LLaVA_Instruct_150k. LLaVA_Instruct_150k contains human-style instruction-response conversations, automatically created by a powerful LLM (OpenAI, 2023) from a generic object detection dataset; see (Liu et al., 2023a) for details. In our paper, we start from a pretrained LLaVA model and only perform a single-stage finetune, updating $\theta_{\text{LLM}}$ and $\theta_{\text{MLP}}$.
@@ 4 Visuomotor instruction tuning
We leverage large Vision Language Models (VLMs) as a generalist to address diverse visuomotor control challenges. Specifically, we transform a typical behavior cloning dataset into an instruction tuning dataset and subsequently finetune a pretrained VLM on this tailored dataset, a process we call Visuomotor Instruction Tuning. The resulting LLaRA framework benefits from the broad, inherent knowledge embedded within the VLM, enabling better visuomotor task learning.
@@
Instruction tuning data from trajectories. Next, we turn a set of expert trajectories $\mathcal{D}$ into a visuomotor instruction tuning dataset. We focus on the stationary robot manipulation scenario featuring a robotic manipulator fixed above a flat table. The visual observation for this setup is captured by a fixed third-person view camera positioned to observe the table surface. For each state transition, we convert the state action pair $(s_{t},a_{t})$ into a single-round conversation. The current visual observation and textual task description, forming $s_{t}$, can be directly assimilated by a VLM as the user instruction. However, the numerical action $a_{t}$ needs conversion into textual format to be generated by a VLM. In contrast to the approach employed by RT-2 (Brohan et al., 2023a), which utilizes special tokens to directly encode numerical action values, we adopt 2D Image Coordinates—normalized to $[0,1]$—to represent the positions in actions. Additionally, any rotational component of the action, denoted as $a_{r}$ (such as the angle of a joint), is expressed textually as ‘‘rotate <r>$a_{r}$</r> degrees’’. To further improve the performance, we incorporate the history of executed actions into the query and structure the response to predict multiple future action steps until the episode concludes. More details are covered in Appendix A.1.
@@
Aggregate multiple images for a single-image VLM. The VLM policy we use here closely follows the design of LLaVA-1.5 (Liu et al., 2024b). The model consumes a single image and language instruction and generates language output. In scenarios where a single observation $s_{t}$ comprises multiple images, systems such as LLaVA suffer from low performance because it was never trained on multiple images (see Tab. 11). Instead, we convert each additional image into a language description with the help of object detection on the additional images (e.g., images in the task description, rather than the main visual observation). The main image is still directly used as the visual input by the VLM without any changes. Compared to inBC, we name this method that takes additional object detection results as Description-Instruct-BC ($\mathcal{D}$-inBC), and our model trained on this data is referred to with the same name as earlier. Examples are also available in Fig. 3, Tab. 5 and Tab. 6.
@@
Figure 3: Example data for visuomotor instruction tuning. (left) BC is the original behavior cloning data (action omitted). inBC is our instruction tuning data created from BC. (right) $\mathcal{D}$-inBC leverages object detection results on the additional images from the task description. Both inBC and $\mathcal{D}$-inBC belong to the LLaRA family.
@@ 5 Supercharging visuomotor instruction tuning dataset
More specifically, given an expert trajectory, we generate single-turn conversations that reveal useful information for learning a policy implicitly. In addition to visual observations, we make use of object labels, locations, and geometries (e.g., rotation) together with task descriptions as expert (or oracle) information. We primarily introduce six auxiliary datasets, augmenting existing inBC or $\mathcal{D}$-inBC to better finetune VLMs. In each dataset, we consider a standard instruction, which is further rephrased multiple times using GPT-4 (OpenAI, 2023) to infuse a reasonable diversity. The answers follow simple rule-based templates that format expert information as expected responses for each conversation. The following paragraphs briefly describe each auxiliary dataset. We kindly refer the readers to Fig. 4 and Appendix A.1.3 for formatting details and qualitative samples.
@@
Figure 4: Auxiliary datasets for visuomotor instruction tuning. Given an input trajectory, we make use of expert information (e.g., object detections) to formulate conversations related to auxiliary semantics (left). Note the same template used among different answer types, highlighted with the same color. One detection demo is shown in the right figure. Please refer to Tab. 7 and Tab. 8 for the actual examples of each dataset.
@@ 6.1 Simulation experiments
We train all methods on these three datasets and evaluate them with 3 levels of difficulties following the test protocol (L1 to L3). Due to missing data in the original VIMA dataset, we were not able to evaluate our approach on L4 properly. For each task, we evaluate all the methods with 20 random seeds and report the average success rates of each level.
Figure 5: Performance on three VIMA subsets. x-axis is the number of expert episodes in the training set.
Effectiveness of LLaRA framework. Figure 5 illustrates the performance of selected methods across various dataset scales. All methods generally benefit from increased training data. Key observations include:
  • inBC consistently surpasses the RT-2 Style baseline, and similarly, $\mathcal{D}$-inBC outperforms $\mathcal{D}$-RT-2 Style. This supports the effectiveness of our approach, which utilizes instruction-tuning style data as well as image coordinates. It is easier to benefit from the existing VLMs pretrained in a conversation style. Methods based on RT-2 Style improve when more robot supervision data is available; however, they significantly underperform compared to our methods when data is limited. Also in Appendix B.2, we confirm even with 100% data, RT-2 Style still underperforms our method trained on only 12% data.
  • $\mathcal{D}$-inBC generally excels over inBC owing to its explicit delineation of object locations. This trend is similarly observed between $\mathcal{D}$-RT-2 Style and RT-2 Style.
@@
Comparison to VIMA. We first clarify the difference between our method and VIMA (Jiang et al., 2023) in terms of the inputs. VIMA takes both front and top view images from the environment while ours only takes the front view. We test the most capable model released by VIMA, which is trained on 660k expert trajectories, and the results are shown in Tab. 1. Compared to VIMA (Jiang et al., 2023), our best model not only achieves better performance but also requires less input and is trained on only 12% of the expert trajectories used in VIMA. The full comparison is in Tab. 9.
Table 1: Comparison to VIMA (Jiang et al., 2023). Our best model not only achieves better performance but also requires less input and is trained on only 12% of the data used in VIMA.
@@
Ablation on auxiliary datasets. We next study the effectiveness of auxiliary datasets. Tab. 2 shows the different combinations of the auxiliary datasets we studied. For each setting, we randomly sample the same amount of examples from each auxiliary dataset and combine them with the converted behavior cloning dataset. The ‘*’ after a ‘✓’ means the reference images that appeared only in the task description are not used to generate this dataset, which makes the dataset less diverse. We train the model on both inBC and $\mathcal{D}$-inBC with different auxiliary dataset settings. On VIMA-0.8k, we control the total number of samples from the auxiliary dataset relative to the samples from the converted BC datasets and train all models for 2 epochs. Results in Fig. 6 show that in general, on VIMA-0.8k, the model performs better with more auxiliary data. In Appendix B.3, we further validate that the model with auxiliary data benefits from the data instead of extra optimization steps.
Figure 6: inBC (top) and $\mathcal{D}$-inBC (bottom) with different auxiliary dataset settings. Each model is trained on VIMA-0.8k for 2 epochs. In general, the model performs better with more auxiliary data.

### v3

@@ Abstract
Vision Language Models (VLMs) have recently been leveraged to generate robotic actions, forming Vision-Language-Action (VLA) models. However, directly adapting a pretrained VLM for robotic control remains challenging, particularly when constrained by a limited number of robot demonstrations. In this work, we introduce LLaRA: Large Language and Robotics Assistant, a framework that formulates robot action policy as visuo-textual conversations and enables an efficient transfer of a pretrained VLM into a powerful VLA, motivated by the success of visual instruction tuning in Computer Vision. First, we present an automated pipeline to generate conversation-style instruction tuning data for robots from existing behavior cloning datasets, aligning robotic actions with image pixel coordinates. Further, we enhance this dataset in a self-supervised manner by defining six auxiliary tasks, without requiring any additional action annotations. We show that a VLM finetuned with a limited amount of such datasets can produce meaningful action decisions for robotic control. Through experiments across multiple simulated and real-world tasks, we demonstrate that LLaRA achieves state-of-the-art performance while preserving the generalization capabilities of large language models. The code, datasets, and pretrained models are available at https://github.com/LostXine/LLaRA.
Figure 1: A real-world demonstration of LLaRA solving an unseen task. In this setting, LLaRA converts only eight thousand simulated expert trajectories from VIMA (Jiang et al., 2023) into a strong instruction-tuning data and generates six auxiliary datasets in a self-supervised manner. With such data and about one thousand additional in-domain images, a pretrained Vision Language Model (VLM) is finetuned into a powerful robot policy that can solve many unseen real-world tasks. At inference, the VLM is queried to generate actions in natural language where the robot action is aligned to image pixel coordinates.
@@ 1 Introduction
Several robot learning approaches also leverage pretrained LLMs / VLMs (Zeng et al., 2022; Brohan et al., 2023c; Driess et al., 2023; Vemprala et al., 2024; Kim et al., 2024; Niu et al., 2024; Zheng et al., 2024) with recent work focusing on low-level robot actions (Brohan et al., 2023b; a; Padalkar et al., 2023). The strong performance of such approaches can be attributed to the extensive world knowledge (e.g., understanding of physics, human common-sense) (Yu et al., 2023; Zhao et al., 2024) and powerful reasoning abilities (Creswell & Shanahan, 2022; Liu et al., 2023b) of underlying language models. However, studies on curating and using conversation-style data for instruction tuning in robotics have been rather limited (Zhen et al., 2024).
Motivated by the promising attributes of VLMs, we explore a process, called Visuomotor Instruction Tuning, of adapting a VLM to a robot action policy that can handle diverse visuomotor control challenges. More specifically, we transform a typical behavior cloning (BC) dataset into a conversational dataset, Instruct-BC (inBC) for VLM finetuning. Given a state described in visual-textual modalities, our VLM is trained to generate suitable actions in text, aligned to image pixel coordinates. Such a formulation based on conversation-style instruction-response data enables us to convert a VLM into a robot action policy effortlessly. However, we also note that the effectiveness of an instruction-tuned VLM depends heavily on the quality of data formulation (Liu et al., 2023a; Ranasinghe et al., 2024b; Kuckreja et al., 2024), which is non-trivial to automate (or, scale) for new domains  (Li et al., 2023; Kuckreja et al., 2024; Hong et al., 2023; Chen et al., 2023; Ranasinghe et al., 2024b). Therefore, to strengthen (or, supercharge) such data, we construct auxiliary datasets based on the same BC dataset, in a self-supervised fashion without any extra action label.
Our overall framework termed LLaRA (Large Language and Robotics Assistant) generates visuomotor instruction data that can efficiently and effectively finetune a VLM into a robot action policy. It is motivated by LLaVA (Liu et al., 2023a), which is designed primarily for vision tasks. Similarly, LLaRA offers the entire framework of data generation, model formulation, and training pipeline for VLMs, now specialized for robot learning. Our key contributions are:
@@
    Formulating robot manipulation tasks into instruction-response pairs described in natural language, which enables successful instruction tuning of a VLM as a policy.
@@
    Aligning robot actions to image pixel coordinates, makes the transfer of a pretrained VLM to the robotics domain more efficient, particularly when the training data is limited.
@@
    Identifying and generating auxiliary data that further enhances robot policy learning in a self-supervised manner.
@@ 2 Related work
PALM-E (Driess et al., 2023) utilizes a pretrained VLM for embodied reasoning and planning. However, it assumes access to additional policies capable of performing low-level skills from a limited vocabulary, with the VLM generating natural language that conditions these low-level commands.
SpatialVLM (Chen et al., 2024) focuses on enhancing the spatial reasoning ability of the VLM, resulting in a VLM that serves as a spatial relationship predictor rather than a robot policy. To embed SpatialVLM into a policy, chain-of-thought spatial reasoning is required.
@@
There are concurrent works (Kim et al., 2024; Yuan et al., 2024; Niu et al., 2024; Zheng et al., 2024) which also utilize a LLaVA style VLM as a robot policy. All such approaches employ a VLM as a robot action policy, processing both visual observations and textual instructions as inputs. OpenVLA (Kim et al., 2024) employs specialized tokens to represent quantized actions, a method akin to RT-2 (Brohan et al., 2023a). OpenVLA also incorporates DINOv2 (Oquab et al., 2023) as an additional visual encoder along with SigLIP (Zhai et al., 2023), compared to a single CLIP (Radford et al., 2021) vision encoder used in LLaVA. RoboPoint (Yuan et al., 2024) introduces a novel point-based action space and a scalable data pipeline tailored for spatial affordance prediction. Actions are delineated via points on an RGB image, which are subsequently translated to 3D space using depth data. This approach eliminates the reliance on predefined action primitives (Liang et al., 2023; Singh et al., 2023), external object detectors (Huang et al., 2023b; Liu et al., 2024a), and iterative visual prompting (Nasiriany et al., 2024). LLARVA (Niu et al., 2024), differs by generating both 2D visual traces in image coordinates and corresponding textual actions as outputs, with the former functioning as an auxiliary task.
In contrast, our method describes robot action using image coordinates in natural language, which creates a more intuitive link between vision and action without modifying the existing VLM model architecture. This enables a much more efficient transfer of the pretrained VLM to our robot VLA model. Moreover, all the aforementioned studies lack of comprehensive investigation into the methods for generating self-supervised auxiliary datasets from existing robot data, as well as the implications of integrating such datasets for the training. Our work LLaRA distinctively addresses this gap, marking a critical divergence from the concurrent studies.
@@ 3 Preliminary: visual instruction tuning
Visual Instruction Tuning is a framework to finetune a large Vision Language Model (VLM) using task-specific instructions. The objective is to develop a general-purpose multimodal model capable of solving diverse vision tasks described by language (Huang et al., 2023a). A typical example of Visual Instruction Tuning is LLaVA (Liu et al., 2023a), whose model consists of three components: a large language model $\theta_{\text{LLM}}$, a visual encoder $\theta_{V}$ and an adapter layer $\theta_{\text{MLP}}$ (see Fig. 2).
@@
Consider a conversation data sample that contains a single image $x_{v}$ and user instruction text $x_{l}$ as inputs. During inference, the visual encoder $\theta_{V}$ and adapter layer $\theta_{\text{MLP}}$ process $x_{v}$ successively to produce visual tokens that can be directly concatenated with the textual tokens of $x_{l}$ in language embedding space of $\theta_{\text{LLM}}$. Next, $\theta_{\text{LLM}}$ autoregressively generates new tokens conditioned on both the visual tokens from $x_{v}$ and text tokens from $x_{l}$. Finally, generated tokens are decoded into natural language as the output of the model.
A two-phase training process is executed in Liu et al. (2023a) with a next-token prediction objective over the $\theta_{\text{LLM}}$ output in each phase. First, only $\theta_{\text{MLP}}$ is trained using image-caption pairs (similar to datasets for CLIP (Radford et al., 2021) training). Then, both $\theta_{\text{MLP}}$ and $\theta_{\text{LLM}}$ are finetuned using a specialized instruction tuning dataset termed LLaVA_Instruct_150k. LLaVA_Instruct_150k contains human-style instruction-response conversations, automatically created by a powerful LLM (OpenAI, 2023) from a generic object detection dataset. This instruction tuning process is what we focus on in this paper. We start from a pretrained LLaVA model and only perform a single-stage finetune, updating $\theta_{\text{LLM}}$ and $\theta_{\text{MLP}}$ based on the instruction tuning data formulation we introduce in this paper.
@@ 4 Visuomotor instruction tuning
We leverage large Vision Language Models (VLMs) as a generalist to address diverse visuomotor control challenges. Specifically, we transform a typical behavior cloning dataset into an instruction tuning dataset and subsequently finetune a pretrained VLM on this tailored dataset, a process we call Visuomotor Instruction Tuning. The resulting LLaRA framework benefits from the broad, inherent knowledge embedded within the VLM and the alignment between robot action and image pixel coordinates, enabling more efficient visuomotor policy learning.
@@
Instruction tuning data from trajectories. Next, we turn a set of expert trajectories $\mathcal{D}$ into a visuomotor instruction tuning dataset. For each state transition, we convert the state action pair $(s_{t},a_{t})$ into a single-round conversation. The current visual observation and textual task description, forming $s_{t}$, can be directly assimilated by a VLM as the user instruction. However, the numerical action $a_{t}$ needs conversion into textual format to be generated by a VLM. In contrast to the approach employed by GATO (Reed et al., 2022), RT-2 (Brohan et al., 2023a), and OpenVLA (Kim et al., 2024), which utilizes special tokens to encode quantized action values (binning tokenization), we adopt 2D Image Coordinates—normalized to $[0,1]$—to represent the positions in actions. Additionally, any rotational component of the action, denoted as $a_{r}$ (such as the angle of a joint), is expressed textually as ‘‘rotate <r>$a_{r}$</r> degrees’’. To further improve the performance, we incorporate the history of executed actions into the query and structure the response to predict multiple future action steps until the episode concludes. More details are covered in Appendix A.1.
Figure 3: Example data for visuomotor instruction tuning. (left) inBC is our instruction tuning data created from BC, only taking textual task description. (right) $\mathcal{D}$-inBC is designed for multi-modal task description that considers reference images by taking their corresponding textual description. Both inBC and $\mathcal{D}$-inBC belong to the LLaRA family.
@@
Aggregate multiple images for a single-image VLM. The VLM policy we use here closely follows the design of LLaVA-1.5 (Liu et al., 2024b). The model consumes a single image and language instruction and generates language output. In scenarios where a single observation $s_{t}$ comprises an observation image and multiple reference images, systems such as LLaVA suffer from low performance because it was never trained on multiple images (see Tab. 15). Instead, we convert each additional reference image into a language description with the help of object detection on the additional images (e.g., images in the task description, rather than the main visual observation). The main image is still directly used as the visual input by the VLM without any changes. Compared to inBC, we name this method that takes additional object detection results as Description-Instruct-BC ($\mathcal{D}$-inBC), and our model trained on this data is referred to with the same name as earlier. Examples are also available in Fig. 3, Tab. 8 and Tab. 9. In Appendix A.1.3 we also discuss another variant of $\mathcal{D}$-inBC that uses natural language instead of structured text to describe the reference images: $\mathcal{D}$-inBC (L).
@@ 5 Supercharging visuomotor instruction tuning dataset
More specifically, given an expert trajectory, we generate conversations that reveal useful information for learning a policy implicitly via scripted templates. In addition to visual observations, we make use of automatically generated object labels, locations, and geometries (e.g., rotation) together with task descriptions as expert (or oracle) information. We primarily introduce six auxiliary datasets, augmenting existing inBC or $\mathcal{D}$-inBC to better finetune VLMs. In each dataset, we consider a standard instruction, which is further rephrased multiple times using GPT-4 (OpenAI, 2023) to infuse a reasonable diversity. The answers follow simple rule-based templates that format expert information as expected responses for each conversation. The following paragraphs briefly describe each auxiliary dataset. We kindly refer the readers to Fig. 4 and Appendix A.1.4 for formatting details and qualitative samples.
@@
Figure 4: Auxiliary datasets for visuomotor instruction tuning. Given an input trajectory, we make use of expert information (e.g., object detections) to formulate conversations related to auxiliary semantics. Please refer to Tab. 11 and Tab. 12 for the actual examples of each dataset.
@@ 6.1 Simulation experiments
We train all methods on these three datasets and evaluate them with 3 levels of difficulties following the test protocol (L1 to L3). Due to missing data in the original VIMA dataset, we were not able to evaluate our approach on L4 properly. For each task, we evaluate all the methods with 20 random seeds and report the average success rates of each level and the average of three levels.
Figure 5: Success rates of the models trained on VIMA subsets. The log-scale x-axis shows the number of expert episodes used in the training set. See Appendix E for the underlying numerical data.
Effectiveness of LLaRA framework. Fig. 5 illustrates the performance of selected methods across various dataset scales. All methods generally benefit from increased training data. Key observations include:
  • inBC consistently surpasses the RT-2 Style baseline, and similarly, $\mathcal{D}$-inBC outperforms $\mathcal{D}$-RT-2 Style. This supports the effectiveness of our approach, which utilizes instruction-tuning style data aligned to image pixel coordinates. It is easier to benefit from the existing VLMs pretrained in a conversation style. Methods based on RT-2 Style improve when more robot supervision data is available; however, they significantly underperform compared to our methods when data is limited. Even with 100% data, RT-2 Style still underperforms our method trained on only 12% data (Appendix B.2).
  • $\mathcal{D}$-inBC generally excels over inBC owing to the additional information from reference images. This trend is similarly observed between $\mathcal{D}$-RT-2 Style and RT-2 Style.
@@
Table 1: Comparison to VIMA (Jiang et al., 2023). Our best model not only achieves better performance but also requires less input and is trained on only 12% of the expert trajectories used in VIMA.
@@
Comparison to VIMA. VIMA takes both front and top view images from the environment while ours only takes the front view. We test the most capable model released by VIMA, which is trained on 660k expert trajectories, and the results are shown in Tab. 1. Compared to VIMA, our best model not only achieves better performance but also requires less input and is trained on only 12% of the expert trajectories used in VIMA. The full comparison is in Tab. 13.
Effectiveness of auxiliary datasets. Tab. 2 shows the different combinations of the auxiliary datasets we studied. For each setting, we randomly sample the same amount of examples from each auxiliary dataset and combine them with the converted behavior cloning dataset. The ‘*’ after a ‘✓’ means the reference images that appeared only in the task description are not used to generate this dataset, which makes the dataset less diverse. We train the model on both inBC and $\mathcal{D}$-inBC with different auxiliary dataset settings. On VIMA-0.8k, we control the total number of samples from the auxiliary dataset relative to the samples from the converted BC datasets and train all models for 2 epochs. Results in Fig. 6 show that in general, on VIMA-0.8k, the model performs better with more auxiliary data. Tab. 3 shows that on VIMA-8k, auxiliary datasets can significantly benefit RT-2 Style models as well.
Figure 6: inBC (top) and $\mathcal{D}$-inBC (bottom) with different auxiliary dataset settings. Each model is trained on VIMA-0.8k for 2 epochs. In general, the model performs better with more auxiliary data. See Appendix E for the underlying numerical data.
@@
Table 3: RT-2 Style benefits from auxiliary datasets, all models are trained on VIMA-8k.
Method L1 (%) L2 (%) L3 (%)
RT-2 Style 3.8 3.1 1.7
RT-2 Style + Aux (D) 36.9 35.8 32.1
inBC 57.3 46.2 42.9
inBC + Aux (D) 59.2 58.8 52.1